Update from Jim & Cesar:
-Fire alarm was cleared.
-It was the p5000etch heat exchanger line that blew, so this is down now.
-The gases have been reset, except for the new HBr which seems to have
an exhaust error.
-Oxygen has been a problem to reset. Jim tried 3 times before the
excess flow switch stayed open. It might be worth asking Air Products
-Lab activities were very light (probably because the ASML is down.)
Monday morning, we'll do a run through of what happened and any follow
up tasks in our lab update meeting. Many thanks to Jim and Cesar for
taking care of things on the scene and John for coaching.